Lena Company has provided the following data for its 2019 operations (ignore income taxes):

2019 revenues were $83,000.

2019 expenses were $49,200.

Dividends declared and paid during 2019 totaled $8,300.

Total assets at December 31, 2019 were $191,000.

Total liabilities at December 31, 2019 were $117,000.

Common stock at December 31, 2019 was $33,000.

Which of the following is correct?

Multiple Choice

  • 2019 net income was $25,500. (A)

  • Total stockholders' equity at December 31, 2019 was $250,000. (B)

  • Retained earnings at December 31, 2019 were $41,000. (C)

  • Retained earnings at December 31, 2019 were $25,500 (D)
Answer #1

Net income = Revenues - Expenses = 83,000 - 49,200

= 33,800 (Option A is wrong)

Total stockholders equity = Total assets - Total Liabilities

= 191,000 - 117,000 = 74,000 (Option B is wrong)

Retained earnings = Net income - dividends paid

= 33,800 - 8,300

= 25,500

Option D is the answer